OpenStudy (anonymous):

differentiate F(x)=3x^2 +4x+100.

OpenStudy (anonymous):

Hi.. hal_stirrup... use the formula for the derivative of x^n for this algebraic equation.. such that .. The forumla says, derivative of x^n is n * x^ (n-1) F ' (x)= 3 ( derivative of x^2) + 4 (derivative of x) + (derivative of 100) = 3 (2x) + 4 * 1 + 0 = 6x+4